Two blocks of different mass are connected by a massless rope which goes over a massless, frictionless pulley. The rope is free to move, and both of the blocks hang vertically. What is the magnitude of the tension in the rope?
Answer #1

Assuming that there is no friction of the pulley, we have...


m_1 * g - T = m_1a

T - m_2 * g = m_2a


Now...


T - m_1 * g = -m_1a

T - m_2 * g = m_2a


2T = m_2a - m_1a + g(m_1 + m_2)

T = (m_2a - m_1a + g(m_1 + m_2))/2
Answer #2

Assuming that there is no friction of the pulley, we have...

m_1 * g - T = m_1a
T - m_2 * g = m_2a

Now...

T - m_1 * g = -m_1a
T - m_2 * g = m_2a

2T = m_2a - m_1a + g(m_1 + m_2)
T = (m_2a - m_1a + g(m_1 + m_2))/2